1.A skateboard was reduced in price from &50 to $43.What was the precent decrease?
Answer by nyc_function(2741)

Let p = percent decrease
(50 - 43)/50 = p/100
7/50 = p/100
50p = 7*100
50p = 700
p = 700/50
p = 14
The percent decrease was 14.